# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ IV Gaussian unsteerable superchannels
In this section, we discuss two special type of Gaussian unsteerable superchannels, that is, maximal Gaussian unsteerable superchannels and Gaussian unsteerable superchannels. Here, we say that a Gaussian superchannel is Gaussian unsteerable if it maps any Gaussian unsteerable channels into Gaussian unsteerable channels; and is maximal Gaussian unsteerable if it maps any maximal Gaussian unsteerable channels into maximal Gaussian unsteerable channels.
Denote by ${\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ the set of all maximal Gaussian unsteerable superchannels, that is,
$\begin{array}[]{rl}&{\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\\ =\{&\Phi\in{\mathcal{SGC}}(H_{A}\otimes H_{B}):\\ &\Phi({\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\},\end{array}$
and by ${\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ the set of all Gaussian unsteerable superchannels, that is,
$\begin{array}[]{rl}&{\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =\{&\Phi\in{\mathcal{SGC}}(H_{A}\otimes H_{B}):\\ &\Phi({\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\}.\end{array}$
@@
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any $(m+n)$-mode Gaussian channel. Then $\phi\in{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ if and only if
$\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(0_{2m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$
@@
Suppose that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian superchannel. By Theorem 1, there exist some ${\chi_{1}}\left(\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},0,0\right),{\chi_{2}}\left(A,Y,\nu\right)\in{\mathcal{GC}}(H_{A}\otimes H_{B})$ such that $\Phi\left(\phi\right)={\chi_{2}}\circ\phi\circ\chi_{1}$ for all $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}(H_{A}\otimes H_{B})$; and moreover, $\Phi\left(\phi\left(K,M,\mathbf{d}\right)\right)=\phi^{\prime}\left(K^{\prime},M^{\prime},\mathbf{d}^{\prime}\right)$ with
$K^{\prime}=AK\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},\ \ M^{\prime}=AMA^{\rm T}+Y.$
@@
Take any $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ and any $\rho=\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in{\mathcal{GS}_{\mathcal{US}}(H_{A}\otimes H_{B})}$. Notice that the CM $\Gamma_{\tau}$ of any $(m+n)$-mode unsteerable Gaussian state $\tau$ can be written as $\Gamma_{\tau}=0_{2m}\oplus Q_{\tau}+P_{\tau}$ with some real matrix $Q_{\tau}\geq i{\Omega_{n}}$ and $P_{\tau}\geq 0$ MKIB . Now, if both $\chi_{1}$ and $\chi_{2}$ are maximal unsteerable, then
$K^{\prime}\Gamma_{\rho}(K^{\prime})^{\rm T}+M^{\prime}$
$=$ $A(K\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\Gamma_{\rho}\Sigma_{m+n}E\Sigma_{m+n})K^{\rm T}A^{\rm T}$
$+AMA^{\rm T}+Y$
$=$ $A[K(0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)})K^{\rm T}+M]A^{\rm T}+Y$
$=$ $A(0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)})A^{\rm T}+Y$
@@
$0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)}=\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\Gamma_{\rho}\Sigma_{m+n}E\Sigma_{m+n},$
$0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)}=K\left(0_{2m}\oplus\sigma_{\rho}^{(1)}+Q_{\rho}^{(1)}\right)K^{\rm T}+M,$
the second equation dues to the maximal unsteerability of $\chi_{1}$, the third equation is because $\phi$ is maximal unsteerable, and the last inequality is as $\chi_{2}$ is maximal unsteerable. It follows that $\Phi\left(\phi\right)\in{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$.
@@
$Y+\left(0_{2m}\oplus i\Omega_{n}\right)-A\left(0_{2m}\oplus i\Omega_{n}\right){A^{\rm T}}\geq 0$ (18)
@@
$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}\geq 0.$ (19)
Noting that $\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}=-\left(0_{2m}\oplus i\Omega_{n}\right)$, Ineq.(19) implies
$\left(0_{2m}\oplus i\Omega_{n}\right)$ (20)
$-\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}E\Sigma_{m+n}\geq 0.$
Comparing Ineqs.(18), (20) and (9) gives $\chi_{1},{\chi_{2}}\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$. In this case, for any $\phi\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$, $K^{\prime}$ and $M^{\prime}$ satisfies

### v2

@@ IV Gaussian unsteerable superchannels
In this section, we discuss two special type of Gaussian unsteerable superchannels, that is, maximal Gaussian unsteerable superchannels and Gaussian unsteerable superchannels. Here, we say that a Gaussian superchannel is Gaussian unsteerable from A to B if it maps any Gaussian unsteerable (from A to B) channels into Gaussian unsteerable (from A to B) channels; and is maximal Gaussian unsteerable from A to B if it maps any maximal Gaussian unsteerable (from A to B) channels into maximal Gaussian unsteerable (from A to B) channels.
Denote by ${\mathcal{GSC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ the set of all $(m+n)$-mode maximal Gaussian unsteerable (from A to B) superchannels, that is,
$\begin{array}[]{rl}&{\mathcal{GSC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}\\ =\{&\Phi\in{\mathcal{GSC}}^{(m,n)}:\\ &\Phi({\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)})\subseteq{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}\},\end{array}$
and by ${\mathcal{GSC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ the set of all $(m+n)$-mode Gaussian unsteerable (from A to B) superchannels, that is,
$\begin{array}[]{rl}&{\mathcal{GSC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}\\ =\{&\Phi\in{\mathcal{GSC}}^{(m,n)}:\\ &\Phi({\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)})\subseteq{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}\},\end{array}$
@@
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{\left(m,n\right)}$ is any $(m+n)$-mode Gaussian channel. Then $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ if and only if
$\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(0_{2m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$ (18)
@@
Thus, by Theorem 8, we can give sufficient conditions of Gaussian superchannels becoming maximal unsteerable and unsteerable, respectively.
Theorem 9.
Assume that $\Phi=\Phi\left(A,E,Y,\nu\right)\in{\mathcal{GSC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian superchannel. If either
@@
(1) there exist maximal Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(K_{1},M_{1},\mathbf{d}_{1})$, $\chi_{2}=\chi_{2}(K_{2},M_{2},\mathbf{d}_{2})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}^{(m,n)}$, or
(2) the matrices $A,E$ and $Y$ satisfy
and
for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, then $\Phi\in{\mathcal{GSC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$.
If either
(3) there exist Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(K_{1},M_{1},\mathbf{d}_{1})$, $\chi_{2}=\chi_{2}(K_{2},M_{2},\mathbf{d}_{2})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}^{(m,n)}$, or
(4) the matrices $A,E$ and $Y$ satisfy
$Y+\left(0_{2m}\oplus i\Omega_{n}\right)-A\left(0_{2m}\oplus i\Omega_{n}\right){A^{\rm T}}\geq 0$
and
$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}\geq 0,$
then $\Phi\in{\mathcal{GSC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.
Proofs of Theorems 8 and 9 are given in Appendix.
In addition, there exists maximal Gaussian unsteerable superchannels which are not Gaussian unsteerable superchannels. For example, take a $(1+1)$-mode Gaussian superchannel $\Phi=\Phi\left(A,E,Y,\nu\right)$ with
[…]
and
[…]
By calculations, we find that $A,E,Y$ satisfy
${w^{\dagger}}Yw+\left|{{w^{\dagger}}A\left({{0_{2}}\oplus{\Omega_{1}}}\right){A^{\rm T}}w}\right|\geq\left|{{w^{\dagger}}\left({{0_{2}}\oplus{\Omega_{1}}}\right)w}\right|$}
for any $w\in\mathbb{C}^{4}$, but do not satisfy
$Y-i\left({{0_{2}}\oplus{\Omega_{1}}}\right)+A\left({{0_{2}}\oplus i{\Omega_{1}}}\right){A^{\rm T}}\geq 0.$}
That implies that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{GSC}}^{(1,1)}_{\mathcal{MUS}(A\rightarrow B)}$, but $\Phi\left(A,E,Y,\nu\right)\notin{\mathcal{GSC}}^{(1,1)}_{\mathcal{US}(A\rightarrow B)}$. Thus, the set of Gaussian maximal unsteerable superchannels differs from that of all Gaussian unsteerable superchannels. Nevertheless, both can be identified as free superchannels within the resource theory of steering for channels in CV systems.
V Conclusion
VI Acknowledments
The authors wish to give their thanks to the referees for many helpful comments to improve the original paper.
This work is partially supported by National Natural Science Foundation of China (11901421, 12571138, 12171290), Shanxi Scholarship Council of China (2025-001) and the special fund for Science and Technology Innovation Teams of Shanxi Province (202304051001035).
Appendix
A proof of Theorem 6.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$ is any $(m+n)$-mode Gaussian steering-annihilating channel. Then for any $\rho\in{\mathcal{GS}}^{(m,n)}$ with CM $\Gamma_{\rho}$, $\phi(\rho)$ is unsteerable from A to B or from B to A.
If $\phi(\rho)\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$, by Eqs.(5)-(6), one has
$K\Gamma_{\rho}K^{\rm T}+M\geq\pm i(0_{2m}\oplus\Omega_{n}),$
and so
$\mathbf{w}^{{\dagger}}(K\Gamma_{\rho}K^{\rm T}+M)\mathbf{w}\geq\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w},$
that is,
holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
Note that, it is shown in PMGH that, for any $N$-mode Gaussian state $\rho$ with CM $\Gamma$, it holds that
$\inf\limits_{\Gamma\geq\pm i\Omega_{N}}{\bf w}^{\dagger}\Gamma{\bf w}=|{\bf w}^{\dagger}\Omega_{N}{\bf w}|,\ \ \forall\mathbf{w}\in\mathbb{C}^{2N}.$ (19)
Thus, by Eq.(19), we have
$\begin{array}[]{rl}&\inf\limits_{\Gamma_{\rho}\geq\pm i(\Omega_{m}\oplus\Omega_{n})}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ =&|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ \geq&\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w},\end{array}$
which implies
for all $\mathbf{w}\in\mathbb{C}^{2(m+n)},$ that is, Ineq.(12) holds.
Symmetrically, one can show that, if $\phi(\rho)\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(B\rightarrow A)}$, then Ineq.(13) holds.
On the other hand, if Ineq.(12) holds, then for any $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, by using Eq.(19), one has
$\geq$ $|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
$\geq$ $\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.$
So $\Gamma_{\phi(\rho)}\geq\pm i(0_{2m}\oplus\Omega_{n})$, which means that $\phi(\rho)$ is unsteerable from A to B. Symmetrically, if Ineq.(13) holds, then $\phi(\rho)$ is unsteerable from B to A. It follows that $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$. ∎
A proof of Theorem 7.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{\left(N\right)}$ is any $N$-mode Gaussian channel.
$\left(2\right)\Rightarrow\left(3\right)$: Note that the CM $\Gamma_{{\rho_{\phi}}}$ of $\rho_{\phi}$} has the form X11
[…]
If ${\rho_{\phi}}\in{\mathcal{GS}}^{(N,N)}_{\mathcal{US}(A\rightarrow B)}$, then
$\Gamma_{{\rho_{\phi}}}+(0_{2N}\oplus i{\Omega_{N}})\geq 0.$ (21)
It is well known ZH that a Hermitian matrix
$W=\left(\begin{array}[]{cc}W_{11}&W_{12}\\ W_{12}^{{\dagger}}&W_{22}\end{array}\right)\geq 0$ (24)
$\Leftrightarrow W_{22}\geq 0,\ \ W_{11}-W_{12}W_{22}^{-1}W_{12}^{{\dagger}}\geq 0$
$\Leftrightarrow W_{11}\geq 0,\ \ W_{22}-W_{12}^{{\dagger}}W_{11}^{-1}W_{12}\geq 0.$
So Ineq.(21) implies
Note that $(T+cI)^{-1}=\frac{1}{c}(I-\frac{T}{c})+o(c^{-2})$ S1 , where $o(c^{-2})$ stands for the Landau little-$o$ which will be neglected when taking to $c\to\infty$. So
If ${\rho_{\phi}}\in{\mathcal{GS}}^{(N,N)}_{\mathcal{US}(B\rightarrow A)}$, then
$\Gamma_{{\rho_{\phi}}}+(i\Omega_{N}\oplus 0_{2N})\geq 0.$ (25)
By the equivalence (24) again, Ineq.(25) implies
$\begin{array}[]{rl}&\cosh 2rK{K^{\rm T}}+M+i\Omega_{N}\\ -&\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N})^{-1}\Sigma_{N}{K^{\rm T}}\geq 0.\end{array}$
So
$0\leq\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M+i\Omega_{N}$
- $\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N})^{-1}\Sigma_{N}{K^{\rm T}}]$
$=$ $\mathop{\lim}\limits_{r\to\infty}[\frac{\cosh^{2}(2r)-\sinh^{2}(2r)}{\cosh 2r}K{K^{\rm T}}+M+i\Omega_{N}]$
$=$ $M+i{\Omega_{N}}.$
Therefore, the statement (2) holds.
[…]
$Y+i{\Omega_{N^{\prime}}}\geq 0$}
and
$X+i{\Omega_{N}}-Z\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}{Z^{\rm T}}\geq 0$}.
If the matrices $K$ and $M$ satisfy $M-iK\Omega_{N}K^{\rm T}\geq 0$, one gets
$KX{K^{\rm T}}+M-KZ{\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}}{Z^{\rm T}}{K^{\rm T}}$
$=$ $K\left({X-Z{{\left({Y+i{\Omega_{N^{\prime}}}}\right)}^{-1}}{Z^{\rm T}}}+i{\Omega_{N}}\right){K^{\rm T}}$
$+(M-iK{\Omega_{N}}{K^{\rm T}})\geq 0.$
Also note that $KX{K^{\rm T}}+M\geq 0$ by the denifition of the channel $\phi$. By ZH again, one achieves
${\Gamma_{\left({\phi\otimes I_{H^{\prime}}}\right)(\rho)}}+i(0_{N}\oplus{\Omega_{N^{\prime}}})\geq 0.$
So $(\phi\otimes{I_{H^{\prime}}})\left(\rho\right)$ is unsteerable from A to B, and hence $\phi$ is steering-breaking.
Now, assume that the matrices $K$ and $M$ satisfy the condition $M+i\Omega_{N}\geq 0$. It is shown that in LHQ that, if the matrices $L,V,W\in\mathcal{M}_{n}(\mathbb{C})$ satisfy the condition $L>0$, $W\geq 0$ and $VLV^{{\dagger}}+W$ invertible, then
$V^{{\dagger}}(VLV^{{\dagger}}+W)^{-1}V\leq L^{-1}.$
For $X$, $K$ and $M$, applying the above result, one achieves
${K^{\rm T}}(KX{K^{\rm T}}+(M+i{\Omega_{N}}))^{-1}K\leq X^{-1}.$
and so
$Y-{Z^{\rm T}}{K^{\rm T}}(KX{K^{\rm T}}+M+i{\Omega_{N}})^{-1}KZ$
$\geq$ $Y-{Z^{\rm T}}X^{-1}{Z}\geq 0.$
By the equivalence (24), one obtains
${\Gamma_{\left({\phi\otimes I_{H^{\prime}}}\right)(\rho)}}+i({\Omega_{N}}\oplus 0_{N}^{\prime})\geq 0.$
So $(\phi\otimes{I_{H^{\prime}}})\left(\rho\right)$ is unsteerable from B to A. Hence $\phi$ is steering-breaking. ∎
A proof of Theorem 8.
@@ VI Acknowledments
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{\left(m,n\right)}$ is any $(m+n)$-mode Gaussian channel and $\rho\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ is any $(m+n)$-mode Gaussian unsteerable state with CM $\Gamma_{\rho}$.
On the one hand, if $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$, then $\phi(\rho)\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$. So
$K\Gamma_{\rho}K^{\rm T}+M\geq\pm i(0_{2m}\oplus\Omega_{n}),$
which implies
for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$. By a similar discussion to that of Eq.(19), one can obtain
$\begin{array}[]{rl}&\inf\limits_{\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n})}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ =&|{\bf w}^{\dagger}K(0_{2m}\oplus\Omega_{n})K^{\rm T}{\bf w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ \geq&\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.\end{array}$
Hence
$|\mathbf{w}^{{\dagger}}K(0_{2m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
On the other hand, if (18) holds, as $\Gamma_{\rho}\geq\pm i\left({0_{2m}\oplus{\Omega_{n}}}\right)$, for any $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, we have
$\geq$ $\inf_{\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n})}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
$=$ $|\mathbf{w}^{{\dagger}}K(0_{2m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}$
$\geq$ $|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
$\geq$ $\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.$
This means $\Gamma_{\phi(\rho)}\geq\pm i(0_{2m}\oplus\Omega_{n})$, and so $\phi(\rho)$ is unsteerable from A to B. It follows that $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$. ∎
A proof of Theorem 9.
Suppose that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian superchannel. By X11 , there exist some $\chi_{1}=\chi_{1}(K_{1},M_{1},\mathbf{d}_{1})$, $\chi_{2}=\chi_{2}(K_{2},M_{2},\mathbf{d}_{2})\in{\mathcal{GC}}^{(m,n)}$ such that $\Phi\left(\phi\right)={\chi_{2}}\circ\phi\circ\chi_{1}=\phi^{\prime}(K^{\prime},M^{\prime},\mathbf{d})$ for all $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}^{(m,n)}$. Note that $K^{\prime}$ and $M^{\prime}$ satisfy
$M^{\prime}={K_{2}}K{M_{1}}{K^{\rm T}}{K_{2}}^{\rm T}+{K_{2}}M{K_{2}}^{\rm T}+{M_{2}}$
and
$K^{\prime}=K_{2}{K}K_{1}.$
Also notice that the CM $\Gamma_{\tau}$ of any $(m+n)$-mode unsteerable (from A to B) Gaussian state $\tau$ can be written as
$\Gamma_{\tau}=0_{2m}\oplus Q_{\tau}+P_{\tau}$
with some real matrix $Q_{\tau}\geq i{\Omega_{n}}$ and $P_{\tau}\geq 0$ MKIB .
Now, if both $\chi_{1}$ and $\chi_{2}$ are maximal unsteerable from A to B, then for any $(m+n)$-mode maximal Gaussian unsteerable channel $\phi=\phi\left(K,M,\mathbf{d}\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ and any Gaussian unsteerable state $\rho=\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$, we have
$\Gamma_{\Phi\left(\phi\right)(\rho)}=\Gamma_{{\chi_{2}}\circ\phi\circ{\chi_{1}}(\rho)}$
$=$ ${K}_{2}K{K_{1}}\Gamma_{\rho}{K_{1}}^{\rm T}{K^{\rm T}}{K_{2}}^{\rm T}$
$+{K_{2}}K{M_{1}}{K^{\rm T}}{K_{2}}^{\rm T}+{K_{2}}M{K_{2}}^{\rm T}+{M_{2}}$
$=$ ${K}_{2}[K(0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)})K^{\rm T}+M]{K}_{2}^{\rm T}+M_{2}$
$=$ ${K}_{2}(0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)}){K}_{2}^{\rm T}+M_{2}$
@@
$0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)}=K_{1}\Gamma_{\rho}K_{1}^{\rm T}+M_{1},$
$0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)}=K\left(0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)}\right)K^{\rm T}+M,$
and the second equation, the third equation, the last inequality due to respectively the maximal unsteerability of $\chi_{1}$, $\phi$ and $\chi_{2}$. So $\Phi(\phi)(\rho)$ is unsteerable from A to B. It follows that $\Phi\left(\phi\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$.
If both $\chi_{1}$ and $\chi_{2}$ are Gaussian unsteerable from A to B, that is, they satisfy Ineq.(9):
$M_{2}+\left(0_{2m}\oplus i\Omega_{n}\right)-K_{2}\left(0_{2m}\oplus i\Omega_{n}\right){K_{2}^{\rm T}}\geq 0$
and
$M_{1}+\left(0_{2m}\oplus i\Omega_{n}\right)-K_{1}\left(0_{2m}\oplus i\Omega_{n}\right){K_{1}^{\rm T}}\geq 0,$
then for any $(m+n)$-mode Gaussian unsteerable channel $\phi({K,M,\mathbf{d}})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$, we have
$M+\left(0_{2m}\oplus i\Omega_{n}\right)-K\left(0_{2m}\oplus i\Omega_{n}\right){K^{\rm T}}\geq 0,$
and so
$=$ ${K_{2}}K{M_{1}}{K^{\rm T}}{K_{2}}^{\rm T}+{K_{2}}M{K_{2}}^{\rm T}+{M_{2}}$
$+({0_{2m}}\oplus i{\Omega_{n}})-K_{2}{K}K_{1}(0_{2m}\oplus i{\Omega_{n}})K_{1}^{\rm T}K^{\rm T}{{K_{2}}^{\rm T}}$
$=$ ${K_{2}}K[M_{1}+({0_{2m}}\oplus i{\Omega_{n}})-K_{1}(0_{2m}\oplus i{\Omega_{n}})K_{1}^{\rm T}]K^{\rm T}{K_{2}}^{\rm T}$
$+{K_{2}}[M+({0_{2m}}\oplus i{\Omega_{n}})-K(0_{2m}\oplus i{\Omega_{n}})K^{\rm T}]{K_{2}}^{\rm T}$
$+{M_{2}}+({0_{2m}}\oplus i{\Omega_{n}})-K_{2}(0_{2m}\oplus i{\Omega_{n}}){{K_{2}}^{\rm T}}$
$\geq$ $0.$
Hence $\Phi\left(\phi\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.
@@
$Y+\left(0_{2m}\oplus i\Omega_{n}\right)-A\left(0_{2m}\oplus i\Omega_{n}\right){A^{\rm T}}\geq 0$
@@
$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}\geq 0.$
As $\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}=-\left(0_{2m}\oplus i\Omega_{n}\right)$, the inequality implies
$\begin{array}[]{rl}&\left(0_{2m}\oplus i\Omega_{n}\right)\\ &-\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}E\Sigma_{m+n}\geq 0.\end{array}$
Thus, for any $\phi(K,M,\mathbf{d})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ with $\Phi(\phi)=\phi^{\prime}(K^{\prime},M^{\prime},\mathbf{d}^{\prime})$, $K^{\prime}$ and $M^{\prime}$ satisfy